OpenStudy (anonymous):

factoring question need help. 3x^2+5x+7=0 the answer is x=-5/6plus/minus i squareroot 59/6

OpenStudy (jhannybean):

Use quadratic formula :)\[\large y = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (jhannybean):

\[\large b = 5 \ ,\ a= 3 \ , \ c= 7\]

OpenStudy (anonymous):

\[x=(-5\pm \sqrt{5^2-4(3)(7)})/2(3)\]\[x=-5/6 \pm \sqrt{25-84}/6 \]\[x=\pm i \sqrt{59}/6 -5/6\]

OpenStudy (jhannybean):

Yeah...you're solving for x,not y >.< my bad!

OpenStudy (anonymous):

Also note \[\sqrt{-59}=\sqrt{-1}\sqrt{59}=i \sqrt{59}\]

OpenStudy (jhannybean):

Good job :)